# Daisy-TTS: Simulating Wider Spectrum of Emotions via Prosody Embedding Decomposition

### v1

@@ Abstract
We often verbally express emotions in a multifaceted manner, they may vary in their intensities and may be expressed not just as a single but as a mixture of emotions. This wide spectrum of emotions is well-studied in the structural model of emotions, which represents variety of emotions as derivative products of primary emotions with varying degrees of intensity. In this paper, we propose an emotional text-to-speech design to simulate a wider spectrum of emotions grounded on the structural model. Our proposed design, Daisy-TTS  [^2] We named the model ‘Daisy’ as a reference to the flower-like arrangement of the structural model of emotions. Project Page: https://rendchevi.github.io/daisy-tts/, incorporates a prosody encoder to learn emotionally-separable prosody embedding as a proxy for emotion. This emotion representation allows the model to simulate: (1) Primary emotions, as learned from the training samples, (2) Secondary emotions, as a mixture of primary emotions, (3) Intensity-level, by scaling the emotion embedding, and (4) Emotions polarity, by negating the emotion embedding. Through a series of perceptual evaluations, Daisy-TTS demonstrated overall higher emotional speech naturalness and emotion perceiveability compared to the baseline.
@@ 1 Introduction
Subtle variations in speaking patterns that make up our prosody, allow us to express a wide spectrum of emotions Cowen et al. (2019). In most cases, emotions may not always be expressed in their "pure" states Scherer and Ekman (2014). Emotions we express may vary in their intensity-level, for instance, a higher intensity of sadness might be distinctly expressed as grief. It may even get more intricate as sometimes we express not just a single but a mixture of different emotions Plutchik (1982).
@@
This intricate spectrum of emotions posits an interesting challenge in designing emotional text-to-speech systems. Most works have adapted psychological theories in their design to properly represent and simulate emotional speech, notably with the discrete Ekman (1992) and dimensional Russell (1980) theories of emotion. The former allows for a direct representation of emotions as discrete labels (e.g. anger, joy), but likely incapable of simulating other variety of emotions. The latter does open continuous emotional intensity simulation, but it reduces the emotion representations down into higher-level affect variables of valence, arousal, and dominance.
Another theory, the structural model of emotion Plutchik (1984), represents a more structural view of emotions variety as derivative products of primary emotions. This would allow a wider representation of emotions, yet still underexplored in the emotional TTS design. To the best of our knowledge, it has only been adapted in a few neural emotional TTS, most notably Zhou et al. (2022b); Tang et al. (2023), where they managed to simulate intensity-level and mixture of emotions through rank-based method and emotion embedding conditioning, respectively.
@@
    Our evaluation shows that our proposed design improves the quality of prior work Zhou et al. (2022b) in terms of speech naturalness, emotion perceivability, and simulation capabilities.
@@ 2.1 Structural Model of Emotion
In the structural model of emotion Plutchik (1982), the sheer variety of emotional expressions is characterized as a derivative product of primary emotions. Figure 2 shows a concise visualization of the model in a flower-like arrangement.
@@ 2.2 Emotion Manifestation in Speech
It is intuitive to assume that speech conveys identifiable emotions, but what part of it plays a role? Speech is roughly composed of lexical (what we say) and non-lexical (how we say) components. Though both components do convey emotions Schuller and Schuller (2020), substantial works have shown that prosody, a non-lexical pattern of tune, rhythm, and timbre, plays a main role in conveying emotion in speech Mozziconacci (2002); Cowen et al. (2019).
@@ 2.3 Emotion Modelling in Text-to-Speech
Modelling emotion in a text-to-speech task usually involves designing a conditional model that simulates emotional speech given text and emotion representation. How emotions are represented will affect the characteristics of emotions that can be simulated by the model. In Lee et al. (2017), emotion is represented as discrete labels (e.g. joy, sadness), though this allows the model to simulate primary emotions by explicitly specifying the input emotion label, it hinders the simulation of more complex characteristics such as emotion intensity and a mixture of emotions. To alleviate this, a few recent works adapt the mentioned structural model of emotions in their design. Zhou et al. (2022b) is one of the firsts that introduce the ability to simulate emotion intensity and secondary emotion through a rank-based emotion attribute vector. Tang et al. (2023) represents emotion as a vector embedding extracted from a pretrained speech emotion recognizer, which also allows the simulation of both characteristics by combining the hidden state of embedding.
The aforementioned models provide a solid framework for simulating a variety of emotion characteristics, but most of them haven’t explicitly incorporated prosody modelling in their design, which as stated in Section LABEL:sec:manifest-emo, prosody should have played a main role in conveying emotion in speech. One notable method to model prosody follows Skerry-Ryan et al. (2018); Wang et al. (2018), from which style embeddings are directly learned from speech.  Zaïdi et al. (2022) adopt this method to specifically learn cross-speaker prosody embedding. Though these methods were not explicitly designed to simulate characteristics of emotions, Rijn et al. (2021) shows that the embedding latent space of this family of models is reliably associated with primary emotions.
@@ 3 Learning Emotionally-separable Prosody Embedding
In this section, we discuss our approach to learning emotionally-separable prosody embeddings through our proposed framework, Daisy-TTS. As in Figure 3, the embeddings were learned by encoding non-lexical speech features through a prosody encoder equipped with an emotion classifier and conditioned them to a text-to-speech backbone.
@@ 3.1 Text-to-Speech Backbone
We chose Grad-TTS Popov et al. (2021) as our backbone. Grad-TTS is a non-autoregressive TTS model based on diffusion probabilistic modelling. We chose this model as it is one the most high-quality open-source TTS at the time of this research, which allow us to focus more on the emotion modelling aspect.
In our implementation, we follow the original architecture of Grad-TTS, which is composed of encoder and decoder modules. We apply a slight modification to the encoder for prosody conditioning. The encoder follows the same structure as Kim et al. (2020), consisting of a 1D convolutional prenet, 6 Transformer blocks, and a linear projection layer. The encoder takes in $x$ as direct input, $u$ is conditioned into the module on feature-level with FiLM conditioning Perez et al. (2017), the module outputs $\mu_{\{x,u\}}$, hidden feature of $\{x,u\}$ assumed to be aligned with corresponding speech, $y$. The decoder follows the same U-Net architecture as in Ronneberger et al. (2015); Ho et al. (2020). It synthesizes $y$ by iteratively decoding a Gaussian noise, $z\sim\mathcal{N}(\mu_{\{x,u\}},I)$, parametrized by the encoder’s output.
@@ 3.2 Emotionally-separable Prosody Encoder
To obtain a vector embedding representation of prosody, $u$, we follow the methods from prosody transfer modelling Skerry-Ryan et al. (2018); Wang et al. (2018); Zaïdi et al. (2022). We define $\mathcal{G}(s)$, a prosody encoder model independent of the backbone model. The main objective of $\mathcal{G}(s)$ is to learn an emotionally-separable prosody embedding, $u$, from some non-lexical features, $s$.
@@
To impose emotion separability, we adopt the approach from Ganin and Lempitsky (2015); Zaïdi et al. (2022). We introduce an auxiliary emotion classifier at the end of $\mathcal{G}(s)$, with the task to predict the primary emotion label of $u$ and backpropagate it through a gradient reversal layer Ganin and Lempitsky (2015).
@@ 3.3 Model Training
Training Objective. Both the backbone and the prosody encoder are trained jointly. There are 2 main training objectives, each corresponding to both models’ respective tasks in synthesizing prosody-controlled speech and imposing emotion separability. For the former, we follow the same training objectives in Popov et al. (2021), and for the latter, we follow the cross-entropy loss for the auxiliary classifier in Zaïdi et al. (2022).
Training Dataset. We found ESD (Emotional Speech Dataset) Zhou et al. (2021, 2022a) is suitable for our problem, an open-source (MIT License) emotinal TTS dataset. We select the subset containing 10 English speakers (5 male, 5 female), each with 350 parallel utterances voice acted in neutral and 4 primary emotions: joy, sadness, anger, and surprise. We follow the original training samples split. With data processing protocol follows Popov et al. (2021), where the texts are converted to phonemes and the speech samples are represented as mel-spectrograms with a sample rate set to 22,050 Hz, and mel bands, filter size, window size, and hop length set to 80, 1024, 1024, 256, respectively.
Training Setting. Our model parameters count is 1̃4,800,000. We trained the model in a single GPU (NVIDIA A100). AdamW Loshchilov and Hutter (2019) is used as the optimizer, with $\beta_{1}=0.9$, $\beta_{2}=0.99$, and constant learning rate of $1e^{-4}$. The model was trained for 310,000 steps with a batch size of 32.
Vocoder Setting. As most neural TTS that use mel-spectrogram as the output, we need a vocoder to decode the synthesized speech. We use the pretrained HifiGAN Kong et al. (2020) vocoder available in the official repository of Grad-TTS  [^1] https://github.com/huawei-noah/Speech-Backbones/tree/main/Grad-TTS, and finetuned it with ESD dataset.
@@ 4.1 Emotion Separability
We examine the emotion separability of u by applying the dimensionality reduction method, t-SNE Van der Maaten and Hinton (2008). Figure 1 shows how each $\textbf{u}^{\varepsilon}$ is well separated from each other. This implies that the encoded information describing primary emotion in u are distinguishable from one another, which conforms to the first emotion characteristics described in Section 2.1.
@@ 4.2 Emotion from Prosody Decomposition
As we want to simulate a range of emotions through $u$, it is important to have controllability over $u$ beyond providing an exemplar speech to the model. Our approach is inspired by Blanz and Vetter (2023). We decompose $u$ as a linear combination of its "prototypes" weighted by parameter vector w:
@@
the above decomposition can be obtained by performing PCA Wold et al. (1987) to the embeddings, with $\overline{u}$ as the mean of embeddings, $n$ as the chosen number of principal components, $v_{i}\in\mathbb{R}^{n\times d}$ as the embedding prototypes in the form of its top-$n$ eigenvectors, and w as the parameter vector associated with $u$.
This type of decomposition assumed that w follow a multivariate Gaussian distribution Egger et al. (2020), with probability density function described as:
@@ 4.4 Simulating Secondary Emotions
A secondary emotion is defined as a mixture between 2 primary emotions Plutchik (1982). With our prosody decomposition, we can interpret a secondary emotion as a mixture between two different $\textbf{w}^{\varepsilon}$, denoted as $\textbf{w}^{\varepsilon_{s}}$. As w follow multivariate Gaussian distribution, we can define the mean and covariance for $\textbf{w}^{\varepsilon_{s}}$ as:
@@ 4.6 Emotions Polarity
Polarity in emotion is defined as the maximum dissimilarity of emotion Plutchik (1982). We found that by negating $\textbf{w}^{\varepsilon}$, we can simulate an emotion that is perceivably the opposite of that emotion. For example, sampling $-\textbf{w}^{\varepsilon_{joy}}$ gives embedding that is similar to sampling from $\textbf{w}^{\varepsilon_{sadness}}$ and vice versa.
@@ 5.1 Evaluation Metrics
Speech Naturalness (MOS) $\uparrow$
Anger Joy Sadness Surprise Bittersweet Delight Pride Disappointment Envy Outrage
Ground Truth 3.844 3.867 3.567 3.789 - - - - - -
Baseline 3.356 3.167 3.200 2.822 3.433 3.178 3.356 3.256 3.478 3.189
Daisy-TTS 3.689 3.844 3.456 3.511 3.667 3.811 3.656 3.911 3.900 3.844
@@
Anger Joy Sadness Surprise Bittersweet Delight Pride Disappointment Envy Outrage
Ground Truth 0.656 0.478 0.556 0.489 - - - - - -
Baseline 0.300 0.233 0.533 0.144 0.211 0.311 0.133 0.177 0.255 0.188
Daisy-TTS 0.533 0.333 0.478 0.300 0.288 0.355 0.166 0.100 0.277 0.233
Table 1: Results of MOS and emotion perception tests on primary and secondary emotions. Daisy-TTS overall achieves higher speech naturalness and emotion perceivability compared to the baseline.
@@
We adopt the tests for emotional speech evaluation conducted in Zhou et al. (2022b); Cowen et al. (2019). Emotional speech samples will be evaluated in terms of its speech naturalness and emotion perceivability. Speech naturalness is evaluated through the MOS (Mean Opinion Score) test, where human annotators were asked to rate the naturalness of the given speech samples on a 5-point Likert scale: Excellent, Good, Fair, Poor, and Bad. Whereas emotion perceivability is evaluated through an emotion perception test, where the annotators were asked to select one or two emotions they perceived.
To accommodate the test, we crowd-sourced our evaluation to human annotators. The details of the crowd-sourcing setup can be found in A.2
@@ 5.2 Baseline Setup
As discussed in Section 2.3, there are mainly 2 candidates for our baselines,  Zhou et al. (2022b) and Tang et al. (2023). We can only utilize Zhou et al. (2022b) as our baseline as the latter is not open-sourced.
Through their rank-based method, the baseline model can simulate primary, secondary, and intensity-level of emotions. The model is a sequence-to-sequence TTS akin to the recognition-synthesis model Zhang et al. (2020). We use the official pretrained model shared by the authors  [^2] https://github.com/KunZhou9646/Mixed_Emotions.

### v2

@@ Abstract
We often verbally express emotions in a multifaceted manner, they may vary in their intensities and may be expressed not just as a single but as a mixture of emotions. This wide spectrum of emotions is well-studied in the structural model of emotions, which represents variety of emotions as derivative products of primary emotions with varying degrees of intensity. In this paper, we propose an emotional text-to-speech design to simulate a wider spectrum of emotions grounded on the structural model. Our proposed design, Daisy-TTS  [^2] We named the model ‘Daisy’ as a reference to the flower-like arrangement of the structural model of emotions. Interactive Demo: https://rendchevi.github.io/daisy-tts/, incorporates a prosody encoder to learn emotionally-separable prosody embedding as a proxy for emotion. This emotion representation allows the model to simulate: (1) Primary emotions, as learned from the training samples, (2) Secondary emotions, as a mixture of primary emotions, (3) Intensity-level, by scaling the emotion embedding, and (4) Emotions polarity, by negating the emotion embedding. Through a series of perceptual evaluations, Daisy-TTS demonstrated overall higher emotional speech naturalness and emotion perceiveability compared to the baseline.
@@ 1 Introduction
Subtle variations in speaking patterns that make up our prosody, allow us to express a wide spectrum of emotions (Cowen et al., 2019). In most cases, emotions may not always be expressed in their "pure" states (Scherer and Ekman, 2014). Emotions we express may vary in their intensity-level, for instance, a higher intensity of sadness might be distinctly expressed as grief. It may even get more intricate as sometimes we express not just a single but a mixture of different emotions (Plutchik, 1982).
@@
This intricate spectrum of emotions posits an interesting challenge in designing emotional text-to-speech systems. Most works have adapted psychological theories in their design to properly represent and simulate emotional speech, notably with the discrete (Ekman, 1992) and dimensional (Russell, 1980) theories of emotion. The former allows for a direct representation of emotions as discrete labels (e.g. anger, joy), but likely incapable of simulating other variety of emotions. The latter does open continuous emotional intensity simulation, but it reduces the emotion representations down into higher-level affect variables of valence, arousal, and dominance.
Another theory, the structural model of emotion (Plutchik, 1984), represents a more structural view of emotions variety as derivative products of primary emotions. This would allow a wider representation of emotions, yet still underexplored in the emotional TTS design. To the best of our knowledge, it has only been adapted in a few neural emotional TTS, most notably (Zhou et al., 2022b; Tang et al., 2023), where they managed to simulate intensity-level and mixture of emotions through rank-based method and emotion embedding conditioning, respectively.
@@
    Our evaluation shows that our proposed design improves the quality of prior work (Zhou et al., 2022b) in terms of speech naturalness, emotion perceivability, and simulation capabilities.
@@ 2.1 Structural Model of Emotion
In the structural model of emotion (Plutchik, 1982), the sheer variety of emotional expressions is characterized as a derivative product of primary emotions. Figure 2 shows a concise visualization of the model in a flower-like arrangement.
@@ 2.2 Emotion Manifestation in Speech
It is intuitive to assume that speech conveys identifiable emotions, but what part of it plays a role? Speech is roughly composed of lexical (what we say) and non-lexical (how we say) components. Though both components do convey emotions (Schuller and Schuller, 2020), substantial works have shown that prosody, a non-lexical pattern of tune, rhythm, and timbre, plays a main role in conveying emotion in speech (Mozziconacci, 2002; Cowen et al., 2019).
@@ 2.3 Emotion Modelling in Text-to-Speech
Modelling emotion in a text-to-speech task usually involves designing a conditional model that simulates emotional speech given text and emotion representation. How emotions are represented will affect the characteristics of emotions that can be simulated by the model. In (Lee et al., 2017), emotion is represented as discrete labels (e.g. joy, sadness), though this allows the model to simulate primary emotions by explicitly specifying the input emotion label, it hinders the simulation of more complex characteristics such as emotion intensity and a mixture of emotions. To alleviate this, a few recent works adapt the mentioned structural model of emotions in their design. (Zhou et al., 2022b) is one of the firsts that introduce the ability to simulate emotion intensity and secondary emotion through a rank-based emotion attribute vector. (Tang et al., 2023) represents emotion as a vector embedding extracted from a pretrained speech emotion recognizer, which also allows the simulation of both characteristics by combining the hidden state of embedding.
The aforementioned models provide a solid framework for simulating a variety of emotion characteristics, but most of them haven’t explicitly incorporated prosody modelling in their design, which as stated in Section 2.2, prosody should have played a main role in conveying emotion in speech. One notable method to model prosody follows (Skerry-Ryan et al., 2018; Wang et al., 2018), from which style embeddings are directly learned from speech.  (Zaïdi et al., 2022) adopt this method to specifically learn cross-speaker prosody embedding. Though these methods were not explicitly designed to simulate characteristics of emotions, (Rijn et al., 2021) shows that the embedding latent space of this family of models is reliably associated with primary emotions.
@@ 3 Learning Emotionally-separable Prosody Embedding
In this section, we discuss our approach to learning emotionally-separable prosody embeddings through our proposed framework, Daisy-TTS. As in Figure 3, the embeddings were learned by encoding non-lexical speech features through a prosody encoder equipped with an emotion discriminator and conditioned them to a text-to-speech backbone.
@@ 3.1 Text-to-Speech Backbone
We chose Grad-TTS (Popov et al., 2021) as our backbone. Grad-TTS is a non-autoregressive TTS model based on diffusion probabilistic modelling. We chose this model as it is one the most high-quality open-source TTS at the time of this research, which allow us to focus more on the emotion modelling aspect.
In our implementation, we follow the original architecture of Grad-TTS, which is composed of encoder and decoder modules. We apply a slight modification to the encoder for prosody conditioning. The encoder follows the same structure as (Kim et al., 2020), consisting of a 1D convolutional prenet, 6 Transformer blocks, and a linear projection layer. The encoder takes in $x$ as direct input, $u$ is conditioned into the module on feature-level with FiLM conditioning (Perez et al., 2017), the module outputs $\mu_{\{x,u\}}$, hidden feature of $\{x,u\}$ assumed to be aligned with corresponding speech, $y$. The decoder follows the same U-Net architecture as in (Ronneberger et al., 2015; Ho et al., 2020). It synthesizes $y$ by iteratively decoding a Gaussian noise, $z\sim\mathcal{N}(\mu_{\{x,u\}},I)$, parametrized by the encoder’s output.
@@ 3.2 Emotionally-separable Prosody Encoder
To obtain a vector embedding representation of prosody, $u$, we follow the methods from prosody transfer modelling (Skerry-Ryan et al., 2018; Wang et al., 2018; Zaïdi et al., 2022). We define $\mathcal{G}(s)$, a prosody encoder model independent of the backbone model. The main objective of $\mathcal{G}(s)$ is to learn an emotionally-separable prosody embedding, $u$, from some non-lexical features, $s$.
@@
To impose emotion separability in the prosody embedding space, we introduce an auxiliary emotion discriminator at the end of $\mathcal{G}(s)$. Our empirical findings, as illustrated in Section 6.2, shows that by maximizing the cross entropy between prosody embeddings and their respective emotion labels can effectively impose emotion separability of the prosody embeddings.
@@ 3.3 Model Training
Training Objective. Both the backbone and the prosody encoder are trained jointly. There are 2 main training objectives, each corresponding to both models’ respective tasks in synthesizing prosody-controlled speech and imposing emotion separability. For the former, we follow the same training objectives in (Popov et al., 2021), and for the latter, we maximize the cross-entropy loss between the prosody embeddings and their respective emotion labels.
Training Dataset. We found ESD (Emotional Speech Dataset) (Zhou et al., 2021, 2022a) is suitable for our problem. We select the subset containing 10 English speakers (5 male, 5 female), each with 350 parallel utterances voice acted in neutral and 4 primary emotions: joy, sadness, anger, and surprise. We follow the original training samples split. With data processing protocol follows (Popov et al., 2021), where the texts are converted to phonemes and the speech samples are represented as mel-spectrograms with a sample rate set to 22,050 Hz, and mel bands, filter size, window size, and hop length set to 80, 1024, 1024, 256, respectively.
Training Setting. Our model parameters count is 1̃4,800,000. We trained the model in a single GPU (NVIDIA A100). AdamW (Loshchilov and Hutter, 2019) is used as the optimizer, with $\beta_{1}=0.9$, $\beta_{2}=0.99$, and constant learning rate of $1e^{-4}$. The model was trained for 310,000 steps with a batch size of 32.
Vocoder Setting. As most neural TTS that use mel-spectrogram as the output, we need a vocoder to decode the synthesized speech. We use the pretrained HifiGAN (Kong et al., 2020) vocoder available in the official repository of Grad-TTS  [^1] https://github.com/huawei-noah/Speech-Backbones/tree/main/Grad-TTS, and finetuned it with ESD dataset.
@@ 4.1 Emotion Separability
We examine the emotion separability of u by applying the dimensionality reduction method, t-SNE (Van der Maaten and Hinton, 2008). Figure 1 shows how each $\textbf{u}^{\varepsilon}$ is well separated from each other. This implies that the encoded information describing primary emotion in u are distinguishable from one another, which conforms to the first emotion characteristics described in Section 2.1.
@@ 4.2 Emotion from Prosody Decomposition
As we want to simulate a range of emotions through $u$, it is important to have controllability over $u$ beyond providing an exemplar speech to the model. Our approach is inspired by (Blanz and Vetter, 2023). We decompose $u$ as a linear combination of its "prototypes" weighted by parameter vector w:
@@
the above decomposition can be obtained by performing PCA (Wold et al., 1987) to the embeddings, with $\overline{u}$ as the mean of embeddings, $n$ as the chosen number of principal components, $v_{i}\in\mathbb{R}^{n\times d}$ as the embedding prototypes in the form of its top-$n$ eigenvectors, and w as the parameter vector associated with $u$.
This type of decomposition assumed that w follow a multivariate Gaussian distribution (Egger et al., 2020), with probability density function described as:
@@ 4.4 Simulating Secondary Emotions
A secondary emotion is defined as a mixture between 2 primary emotions (Plutchik, 1982). With our prosody decomposition, we can interpret a secondary emotion as a mixture between two different $\textbf{w}^{\varepsilon}$, denoted as $\textbf{w}^{\varepsilon_{s}}$. As w follow multivariate Gaussian distribution, we can define the mean and covariance for $\textbf{w}^{\varepsilon_{s}}$ as:
@@ 4.6 Emotions Polarity
Polarity in emotion is defined as the maximum dissimilarity of emotion (Plutchik, 1982). We found that by negating $\textbf{w}^{\varepsilon}$, we can simulate an emotion that is perceivably the opposite of that emotion. For example, sampling $-\textbf{w}^{\varepsilon_{joy}}$ gives embedding that is similar to sampling from $\textbf{w}^{\varepsilon_{sadness}}$ and vice versa.
@@ 5.1 Evaluation Metrics
Speech Naturalness (MOS; 95% CI) $\uparrow$
Anger Joy Sadness Surprise Delight Pride Envy
Ground Truth 3.844±0.129 3.867±0.123 3.567±0.133 3.789±0.133 - - -
Baseline 3.356±0.189 3.167±0.201 3.200±0.186 2.822±0.246 3.178±0.209 3.356±0.204 3.478±0.186
Daisy-TTS 3.689±0.192 3.844±0.180 3.456±0.196 3.511±0.186 3.811±0.182 3.656±0.166 3.900±0.172
@@
Anger Joy Sadness Surprise Delight Pride Envy
Ground Truth 0.656 0.478 0.556 0.489 - - -
Baseline 0.300 0.233 0.533 0.144 0.311 0.133 0.255
Daisy-TTS 0.533 0.333 0.478 0.300 0.355 0.166 0.277
Table 1: Results of MOS and emotion perception tests on primary and secondary emotions (continued in Appendix A.1). Daisy-TTS overall achieves higher speech naturalness and emotion perceivability compared to the baseline.
@@
We adopt the tests for emotional speech evaluation conducted in (Zhou et al., 2022b; Cowen et al., 2019). Emotional speech samples will be evaluated in terms of its speech naturalness and emotion perceivability. Speech naturalness is evaluated through the MOS (Mean Opinion Score) test, where human annotators were asked to rate the naturalness of the given speech samples on a 5-point Likert scale: Excellent, Good, Fair, Poor, and Bad. Whereas emotion perceivability is evaluated through an emotion perception test, where the annotators were asked to select one or two emotions they perceived.
To accommodate the test, we crowd-sourced our evaluation to human annotators. The details of the crowd-sourcing setup can be found in A.3
@@ 5.2 Baseline Setup
As discussed in Section 2.3, there are mainly 2 candidates for our baselines,  (Zhou et al., 2022b) and (Tang et al., 2023). We can only utilize (Zhou et al., 2022b) as our baseline as the latter is not open-sourced.
Through their rank-based method, the baseline model can simulate primary, secondary, and intensity-level of emotions. The model is a sequence-to-sequence TTS akin to the recognition-synthesis model (Zhang et al., 2020). We use the official pretrained model shared by the authors  [^2] https://github.com/KunZhou9646/Mixed_Emotions.